### v1

# Analyzing the dense matter equation of states in the light of the compact object HESS J1731-347

Abstract
The recent mass ($0.77\pm^{0.20}_{0.17}M_{\odot}$) and radius ($10.4\pm^{0.86}_{0.78}\text{km}$) measurement of HESS J1731-347 made it one of the most fascinating object if it is indeed a neutron star. In this work, we examine the current status of the dense matter equation of states in the context of this compact object being a neutron star. We use three sets of equation of states corresponding to the three classes - neutron stars, strange stars, and hybrid stars and perform bayesian model selection on them. Our results show that for hadronic models, the EoS is preferred to be stiff at the intermediate densities. This makes the Brueckner-Hartree-Fock approximation and models based on Effective interactions deviate from current astrophysical observations on the inclusion of HESS J1731-347. Furthermore, for the strange star family, the equation of states composed of three flavor quarks prefers relatively smaller bag parameters. Analyzing the hybrid family of equation of states consisting of a first-order phase transition revealed preferences for early first-order phase transition. Comparing all the preferred equations of state among each family it was found that the current astrophysical constraints most prefer the hybrid equation of states.
preprint: APS/123-QED
I Introduction
Quantum chromodynamics (QCD) predicts the existence of quarks and gluons in a deconfined state at higher densities [1, 2] hinting towards a possible phase transition (PT) at some intermediate densities [3]. The dense core of neutron stars (NSs) lies at the lower end of the intermediate densities (typical central densities lie in the range of 2-8 times that of the nuclear saturation density ($n_{s}=0.16\>\text{fm}^{-3}$), making them one of the most fascinating compact objects to study [4]. Terrestrial-based laboratories are yet to reach such densities, further fuelling our interest in these extreme objects. The core of the NSs is still a mystery, with several models suggesting the presence of only hadronic matter [5, 6, 7]. The possibility of PT at intermediate densities opens the possibility of hybrid stars [8, 9, 10, 11, 12], having an inner core composed of quark matter followed by an outer core of hadronic matter [13, 14, 15, 16, 17, 18]. There also exists another unique category of stars called strange stars (SSs) made up of strange quark matter (SQM) [19, 20, 21]. Witten suggested that the SQM is the absolute ground state consisting of $u$,$d$, and $s$ quarks supporting the idea of SSs [22]. The presence of SQM at the core of the star softens the equation of state (EoS) with possible violations in the maximum mass limit of the star imposed by PSR J0348+0432 [23] of mass $2.01\pm 0.04M_{\odot}$.
The only way to demystify the core of NSs is with astrophysical observations. Apart from the mass measurements of PSR J0348+0432, we also have the mass measurement of PSR J0740+6620 [24] which is $2.08\pm 0.7M_{\odot}$. Simultaneous mass and radius measurements from NICER of PSR J0740+6620 [25, 26] and PSR J0030+0451 [27, 28] have helped in constraining the EoSs. The EoSs have been further constrained from the binary tidal deformability measurement ($\tilde{\Lambda}<720$) imposed by the GW170817 event of the binary neutron star merger reported by the LIGO collaboration [29, 30, 31]. These observations have been successful in narrowing down the EoS band constructed from model agnostic approaches [32, 33]. With more improved observations in the future, the EoS band is expected to get thinner.
However, the recent observation of HESS J1731-347 [34] started raising a few eyebrows with a mass and radius measurement of $0.77\pm^{0.20}_{0.17}M_{\odot}$ and $10.4\pm^{0.86}_{0.78}\text{km}$ respectively, making it the lightest known compact star till date. Since such low mass and radius are inherently favored by SSs, we cannot rule out the possibility of the idea of SQM inside the star. Several significant works followed this observation, each of them trying to explain the properties of this star [35, 36, 37, 38, 39]. However, one of the best way to examine any observation is to do a statistical analysis with the data (basically different models) that one has. Bayesian analysis is an important tool that is being used significantly to constrain nuclear models [40, 41, 42]. Recently, Ref [43] showed, using a bayesian framework, the impact of the compact object in the context of nuclear matter.
Bayesian inference techniques have been primarily used for constraining GW observation parameters [44, 45, 46] with Ref [47] developing a model selection technique using various astrophysical observations to compare nuclear matter EoSs. In this paper, we perform bayesian model selection to explore the implications of the compact object HESS J1731-347 along with other astrophysical observations in the light of a NS, a strange star, or a hybrid star. We consider several nuclear matter EoS models, models consisting of three flavored quarks, and also hybrid models based on first-order phase transition (FOPT). Lastly, we compare the three sets of EoSs to analyze which family of EoSs explains the current astrophysical observations the best.
The paper is arranged in the following way. Section II discusses the formalism adopted in this work to describe the EoS models and also the model selection technique and how it is implemented in this work. The results are described in section III, and finally, important results and conclusion are summarized in section IV.
II Formalism
II.1 Hadronic EoSs
For the purpose of hadronic EoSs, we make use of the publicly available EoSs in the CompOSE repository [48]. All these EoSs consist of $npe\mu$ nuclear matter. We have considered 7 EoS models, namely the Density-dependent Relativistic Mean Field (RMF) model, the Effective Interaction model, the Thomas Fermi approach, the RMF model, the Brussels-Montreal energy density functional, non-linear RMF model, and the Brueckner-Hartree-Fock approximation. Along with these, we have also included APR(APR) (an EoS based on variational techniques), CMGO(GDFM-I) (an EoS based on density-dependent covariant density functional), and PT(GRDF2-DD2) (an EoS based on generalized relativistic density functional). More details about the EoSs used can be found in table 1.
EoS $R_{0.77}$(km) $R_{1.4}$(km) $R_{2.0}$(km) $\Lambda_{1.4}$ $M_{max}(\text{M}_{\odot})$
GPPVA(TW)NSunifiedInnerCrust-core [49, 50] 12.75 12.33 11.42 401 2.07
SPG(M2)unifiedNSEoS [51, 52] 12.45 12.63 12.58 518 2.42
SPG(M4)unifiedNSEoS [51, 53] 12.18 12.31 12.22 433 2.35
SPG(M5)unifiedNSEoS [51] 13.15 13.42 13.65 772 2.71
SPG(M3)unifiedNSEoS [51] 12.44 12.65 12.95 523 2.69
GPPVA(DD2)NSunifiedInnerCrust-core [54] 13.06 13.19 13.14 683 2.42
SPG(M1)unifiedNSEoS [51] 12.78 12.8 12.87 534 2.54
RG(SkMp) [55, 56] 12.54 12.5 11.5 467 2.11
RG(SkI4) [55, 57] 12.18 12.38 11.74 458 2.18
RG(SKb) [55, 58] 11.82 12.21 11.69 404 2.2
RG(SLY2) [55, 59] 11.93 11.79 10.7 307 2.06
RG(SK255) [55, 60] 13.56 13.15 12.03 580 2.15
RG(SLY230a) [55, 61] 11.90 11.83 11.05 324 2.11
RG(SKa) [55, 58] 13.02 12.92 12.16 558 2.22
RG(Rs) [55, 62] 13.01 12.93 11.84 553 2.12
VGBCMR(D1MStar) [63] 11.67 11.71 10.47 314 2.00
RG(SkI3) [55, 57] 13.53 13.55 12.80 771 2.25
RG(SK272) [55, 60] 13.61 13.32 12.48 636 2.24
RG(SLY9) [55, 59] 12.53 12.47 11.70 444 2.16
RG(SKI2) [55, 57] 13.51 13.48 12.45 733 2.17
RG(SkI6) [55, 57] 12.33 12.49 11.88 481 2.2
RG(SLY4) [55, 64] 11.84 11.7 10.62 295 2.06
RG(SkI5) [55, 57] 14.10 14.08 13.12 947 2.25
XMLSLZ(DD-LZ1) [6, 65] 12.52 13.15 13.34 732 2.56
XMLSLZ(DDME2) [6, 66] 12.74 13.2 13.22 712 2.48
XMLSLZ(NL3) [6, 67] 14.30 14.59 14.60 1264 2.77
XMLSLZ(PKDD) [6, 68] 13.75 13.63 13.14 765 2.33
XMLSLZ(GM1) [6, 69] 13.47 13.76 13.38 892 2.36
XMLSLZ(TM1) [6, 70] 14.31 14.28 13.51 1045 2.18
XMLSLZ(DDME-X) [6, 71] 12.81 13.37 13.49 792 2.56
XMLSLZ(MTVTC) [6, 72] 13.25 13.1 11.43 600 2.02
XMLSLZ(TW99) [6, 50] 12.35 12.27 11.35 405 2.08
PCGS(PCSB1) [73, 74] 12.98 13.25 12.67 624 2.19
PCGS(PCSB0) [73, 75] 13.04 13.3 13.28 713 2.53
ABHT(QMC-RMF2) [76] 12.00 12.03 11.02 354 2.04
ABHT(QMC-RMF3) [76, 77] 12.33 12.26 11.61 386 2.15
ABHT(QMC-RMF4) [76, 49] 12.00 12.35 12.04 420 2.21
PCP(BSK26) [51, 78] 11.7 11.77 11.18 323 2.17
PCP(BSK25) [51, 79] 11.97 12.37 12.10 476 2.22
PCP(BSK24) [51, 80] 12.26 12.5 12.27 514 2.28
PCP(BSK22) [51, 81] 12.97 13.04 12.58 624 2.26
GPPVA(NL3wrL55)NSunifiedInnerCrust-core [82] 13.32 13.76 14.06 939 2.75
GPPVA(FSU2)NSunifiedInnerCrust-core [83] 14.53 13.29 11.35 816 2.37
GPPVA(FSU2H)NSunifiedInnerCrust-core [84] 12.91 13.29 10.26 750 2.37
GPPVA(TM1e)NSunifiedInnerCrust-core [85] 13.02 13.16 10.59 661 2.12
BL(chiral)withUnifiedCrust [86] 12.60 12.27 11.13 386 2.08
BL(chiral)WithCrust [86, 87] 12.62 12.31 11.13 385 2.08
APR(APR) [88] 11.31 11.33 10.85 248 2.19
CMGO(GDFM-I) [52] 12.72 12.81 12.46 533 2.31
PT(GRDF2-DD2)coldNS [89, 54] 12.84 13.17 13.07 686 2.42
Table 1: List of all the 50 hadronic EoS with the corresponding radius of a $0.77M_{\odot}$, $1.4M_{\odot}$, and $2M_{\odot}$ star, respectively; along with the Tidal Deformability measurement of the $1.4M_{\odot}$ star and the maximum mass given by each EoS. EoSs in red are based on the density-dependent relativistic mean field (RMF) model, ones in crimson are based on Effective interactions, ones in brown are based on Thomas Fermi approach, ones in darkgreen are based on RMF approximations, ones in violet are based on Brussels-Montreal energy density functional, ones in magenta are based on non-linear RMF models and the ones in peach are based on Brueckner-Hartree-Fock approximation. The last three equations are based on variational techniques, density-dependent covariant density functional, and generalized relativistic density functional, respectively. The information of the rejected hadronic EoSs is presented with bold text.
II.2 Quark Matter EoSs
[…]
II.3 EoSs with First-Order PT
Assuming the phase transition to be of FOPT we construct the hybrid EoS from the hadronic and modified MIT bag-model quark matter EoS. The jump/transition from the hadronic phase (HP) to the quark phase (QP) happens at a particular pressure when the chemical potential of the quark phase becomes less than the chemical potential of the hadronic phase. Although this happens at a particular pressure and chemical potential, and they remain smooth throughout, there is a discontinuity in the energy density (and density) corresponding to the latent heat required for the transition. The transition is modeled using Maxwells construction as $p_{\rm HP}(\mu_{\rm c})=p_{\rm QP}(\mu_{\rm c})$ where $\mu_{\rm c}$ is the critical baryonic chemical potential where the transition occurs.
II.4 Bayesian Model Selection
We adopt a bayesian model selection approach to compare various models of EoS. Each unique EoS is considered a model, and we use Bayes’ theorem, defined as:
$P(M|d,I)=\dfrac{P(d|M,I)P(M|I)}{P(d|I)}$ (1)
where $M$ refers to a model (EoS), $I$ refers to any background information that we have, and d refers to the astrophysical data. $P(M|d,I)$ is the posterior probability of the model, $P(d|M,I)$ is the marginalized-likelihood (evidence) for the data, $P(M|I)$ is the prior probability and $P(d|I)$ is a constant term.
The evidence value, $P(d|M,I)$, can be obtained by marginalizing over the parameters of the model as :
[…]
where $\theta$ refers to the parameters of the model, $P(d|\theta,M,I)$ is the likelihood function of the parameters, and $P(\theta|M,I)$ is a prior probability on the parameters given the information of the model. The evidence value for a model is independent of other models and remains constant irrespective of the number of models evaluated simultaneously.
In order to compare two different models ($M_{1}$ and $M_{2}$), we find out the odds ratio between them, which is defined as:
$\mathcal{O}_{M_{1}}^{M_{2}}=\dfrac{P(M_{2}|d,I)}{P(M_{1}|d,I)}=\dfrac{P(d|M_{2},I)}{P(d|M_{2},I)}\times\frac{P(M_{2}|I)}{P(M_{1}|I)}$ (3)
For uniformity, we can take the ratio of the priors to be equal to one so that $P(M_{2}|I)=P(M_{1}|I)$. By doing so, we avoid the preference for one model over the other. This choice can be modified depending on the background information.
Hence, the odds ratio is redefined as:
$\mathcal{O}_{M_{1}}^{M_{2}}=\dfrac{P(d|M_{2},I)}{P(d|M_{1},I)}$ (4)
If the value of the odds ratio is much greater than 1, then model $M_{2}$ is preferred over model $M_{1}$. If the ratio is much smaller than 1, then the inverse is true.
If we perform the analysis for multiple datasets, such that $d=\{d_{k}\}$, then:
$P(\{d_{k}\}|M,I)=\prod_{k}P(d_{k}|M,I)$ (5)
The odds ratio then finally takes the form:
$\mathcal{O}_{M_{1}}^{M_{2}}=\prod_{k}\frac{P(d_{k}|M_{2},I)}{P(d_{k}|M_{1},I)}$ (6)
For our analysis, $d=\{d_{\rm GW},\ d_{\rm HESS},\ d_{\rm NICER}\}$ refers to the three sets of astrophysical observations we have used. The mass and tidal deformability ($\Lambda$) measurements from GW170817 [29, 30, 93] serves as $d_{\rm GW}$. For $d_{\rm NICER}$, the mass and radius measurements from PSR J0030+0451 [27] and PSR J0740+6620 [25] serve as input. Similarly, the mass and radius measurements from HESS J1731-347 [34] form $d_{\rm HESS}$. As both $d_{\rm HESS}$ and $d_{\rm NICER}$ consist only of the mass and radius measurements, the evidence calculation is the same for them.
First, let us calculate the evidence for the mass-radius measurements for the PSR J0030+0451, PSR J0740+6620, and HESS J1731-347 observations. Since we consider each EoS as a model, we replace ‘$M$’ with ‘EoS’ in eq (2). For every EoS, we solve the Tolman-Oppenheimer-Volkoff (TOV) equations [94] to obtain the mass-radius curve, also known as the MR curve. Our EoS can be parametrized either by using the mass or radius values obtained after solving the TOV equations. In our scenario, we use the mass values, and hence the evidence for the NICER observations is given by:
$P(d_{\rm NICER}|EoS,I)$ $=\>\int_{m_{min}}^{m_{max}}P(d_{\rm NICER}|m,$
$R(m,EoS),EoS,I)$
×$\times P(m|EoS,I)dm$ (7)
where $P(m|EoS,I)$ is the prior distribution on our parameter and $P(d_{\rm NICER}|m,R,EoS,I)$ is the likelihood of the data.
Similarly, for the HESS observation, the evidence is given as :
$P(d_{\rm HESS}|EoS,I)$ $=\>\int_{m_{min}}^{m_{max}}P(d_{\rm HESS}|m,$
$R(m,EoS),EoS,I)$
×$\times P(m|EoS,I)dm$ (8)
with $P(d_{\rm HESS}|m,R,EoS,I)$ being the likelihood of the data.
Without loss of generality, we can choose a uniform prior on mass [47, 43]. It is given by:
$\begin{split}P(m|EoS,I)&=\frac{1}{m_{max}-m_{min}};\>\text{$m_{min}\leq m\leq m_{max}$}\\ &=0\hskip 62.59596pt;\text{everywhere else}\end{split}$ (9)
$m_{max}$ is the maximum mass of the EoS obtained after solving the TOV equation. We fix $m_{min}$ equal to $0.5M_{\odot}$. To construct the likelihoods $P(d_{\rm NICER}|m,R,EoS,I)$ and $P(d_{\rm HESS}|m,R,EoS,I)$, we use a Gaussian kernel density estimation (KDE) with the mass and radius samples from NICER and HESS.
To calculate the evidence for the GW data, we parameterize the two masses of binaries ($m_{1}$,$m_{2}$) and their corresponding tidal deformabilities ($\lambda_{1}$,$\lambda_{2}$) as:
$P(d_{\rm GW}|EoS,I)$ $=\int_{m_{2}}^{M_{max}}dm_{1}\int_{M_{min}}^{m_{1}}P(d_{\rm GW}|m_{1},m_{2},$
$\lambda_{1}(EoS,m_{1}),\lambda_{2}(EoS,m_{2}),EoS,I)$
×$\times P(m_{1},m_{2}|EoS,I)dm_{2}$ (10)
To solve eq (10) we make use of the chirp mass [95] given by:
$\mathcal{M}_{chirp}=\dfrac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}=1.186M_{\odot}$ (11)
Where $m_{1}$ and $m_{2}$ are the masses of the primary and secondary neutron stars having a mass ratio, $q={m_{2}}/{m_{1}}\geq 0.73$ inferred from GW170817 observation [93]. Doing so reduces the parameters needed to evaluate the integral for the evidence of GW170817. We also use the same prior distribution as eq (9) for the GW observation. We construct the likelihood using a multivariate Gaussian KDE with the mass and tidal deformability samples from the observation.
All of the evidence integrals were performed using PyMultinest, which is a Python package for implementing the MultiNest algorithm. It offers efficient evidence calculation for multi-modal data. Furthermore, all likelihood distributions were constructed using the multivariate KDE method of Statsmodels.
III Results
For our analysis, the observations we have used are $(i)$ GW170817, $(ii)$ Three X-Ray sources, namely PSR J0030+0451[27], PSR J0740+6620 [25], and HESS J1731-347 [34].
For the three sets of EoSs that we have considered in our analysis, we have 50 hadronic EoSs, 58 strange matter EoSs, and 637 hybrid EoSs. Since the hybrid EoSs were constructed using Maxwell construction, for each nuclear EoS considered, we agnostically generated a family of hybrid EoSs, resulting in a large number of hybrid EoSs. We adopt Jeffrey’s [96] scale for the log of the odds Ratio values. Note that Jeffrey’s scale is defined for the Bayes Factor and not the odds ratio, but upon taking the ratio of the prior of each model to be unity, the odds ratio becomes equal to the Bayes Factor. According to the scale, if $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-0.5, 0), then although there is evidence for model $M_{1}$, it is not worth more than a bare mention. If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-1, -0.5), then there is ‘substantial’ evidence for (against) model $M_{1}$ ($M_{2})$. If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-2, -1), then there is ‘strong’ evidence for (against) model $M_{1}$ ($M_{2}$). If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ is smaller than -2, then there is ‘decisive’ evidence for (against) model $M_{1}$ ($M_{2}$), and we can reject model $M_{2}$. Utilizing this, we present our analysis in the next subsections.
III.1 Hadronic EoS
Figure 1: Odds ratio plot of SPG(M3) with other hadronic EoSs. Following Jeffrey’s scale, the region between $(-0.5,0)$ (shaded with ‘plum’) is the region in which, if a model lies, it provides evidence for the base model, but it is not worth more than a bare mention (indecisive). The region between $(-1,-0.5)$ (shaded with ‘lightcoral’) is the region of substantial evidence for the base model, and the region between $(-2,-1)$ (shaded with ‘lightpink’) is the region of strong evidence for the base model. The region beyond $-2$ (shaded with ‘mistyrose’) is the region of decisive evidence for the base model. The histogram for the odds ratio value of an EoS is depicted in the corresponding color of the respective region it lies in; for example, if an EoS is situated in the decisive region, its histogram is colored in ‘mistyrose’.
After evaluating the evidence value for each EoS, the EoS with the highest evidence value is SPG(M3) (with unified NS EoS), which uses density-dependent relativistic mean-field approximation. fig. 1 shows the odds ratio plot of each EoS with respect to SPG(M3).
Upon utilizing Jeffrey’s scale, as discussed before, there are exactly 23 EoS that can be decisively rejected based on their odds ratio value with respect to SPG(M3). table 1 highlights the rejected EoSs with bold text. Additionally, there are 9 EoSs with odds ratio values situated in the ‘indecisive’ region (see fig. 1 for reference). They are based on the following EoS models: density-dependent RMF model, Thomas Fermi approximation, RMF model, and generalized relativistic density functional.
Figure 2: Left: EoS plot of the hadronic EoSs. The grey patch shows the region where we see a change in the stiffness of EoS. The fig. 3 shows the speed of analysis in the region of grey patch. Right: Mass-Radius curves of the hadronic EoSs. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
Using table 1 as a reference, it can be observed that except the EoSs based on Brussels-Montreal energy density functional, every other type of EoS has at least one EoS that is being rejected. However, none of the EoSs based on the Brussels-Montreal energy density functional lie in the ‘indecisive’ region. On the contrary, while one EoS based on the density-dependent RMF model can be decisively rejected, 5 (out of 7) of the EoSs lie in the ‘indecisive’ region, with SPG(M3) as the best performing EoS. This analysis suggests that the density-dependent RMF model best satisfies all the current astrophysical observations. Furthermore, it should be noted that all EoSs based on the Brueckner-Hartree-Fock approximation were rejected, leading us to the conclusion that it is the least plausible EoS type that could explain the current observations, including HESS J1731-347.
It was found that 12 (out of 16) Effective interactions based EoSs could be decisively rejected. Among the remaining four EoSs, none are situated in the comparable region; hence, among the EoS types that could not be decisively rejected, the Effective interaction type EoS is the least effective model in our analysis. Furthermore, APR(APR) (based on variational techniques) lies in the ‘strong’ region, CMGO(GDFM-1) (based on density-dependent covariant density functional) lies in the ‘strong’ region, and PT(GRDF2-DD2) (based on generalized relativistic density functional) lies in the ‘substantial’ region. In fig. 2, we show the comparison of the EoSs that were rejected along with their mass-radius curves. We see that at intermediate density ranges, the EoSs which are softer (having less speed of sound) are rejected compared to the stiffer EoSs, which is also shown in fig. 3.
The adiabatic speed of sound ($c_{s}=\sqrt{\partial p/\partial\epsilon}$) is an important quantity as it determines the slope of the EoS. A stiffer EoS would tend to explain higher-mass stars. fig. 3 shows an increasing behaviour in the speed of sound which is characteristic of hadronic matter EoSs [97, 98, 99, 100, 101]. Contrary to the scenario when we do not consider the observation of HESS J1731-347 none of the hadronic EoSs used in this work could be decisively rejected as shown in fig. 4. The rejection of the EoSs from our analysis shows that all the current astrophysical observations prefer a stiffer hadronic EoS at intermediate densities.
Figure 3: The speed of sound corresponding to the shaded region in fig. 2 (left) is shown here. The y-axis represents the speed of sound in geometrized units, and the x-axis represents the energy density. The color nomenclature remains the same as previous plots.
Figure 4: Mass-Radius curves of the hadronic matter EoSs. The figure illustrates the rejected and accepted hadronic EoSs when HESS J1731-347 is not considered. Since none of the EoSs are rejected, all EoSs are plotted in salmon.
III.2 Strange EoS
Figure 5: Odds ratio plot of the strange matter EoS with $x_{v}=0.6$ and B = 139.5 with other strange matter EoSs. The color scheme is the same as in fig. 1. The ticks on the y-axis refer to strange matter EoSs with the corresponding $x_{v}$ and B parameters.
The strange matter EoSs were constructed by varying the vector coupling constant ‘$x_{v}$’ and bag parameters ‘B’ by small intervals as discussed in section II.2. Out of the 58 constructed EoSs, the EoS with $x_{v}=0.6$ and B=139.5 possesses the highest evidence value.
Figure 6: Left: EoS plot of the strange matter EoSs; Right: Mass-Radius curves of the strange matter EoSs. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
Figure 7: Mass-Radius curves of the strange matter EoSs. The figure illustrates the rejected and accepted strange matter EoSs when HESS J1731-347 is not considered. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
The fig. 5 shows the odds ratio plot of each quark matter EoS in comparison to the EoS with $x_{v}=0.6$ and B = 139.5. There are 12 EoSs situated in the ‘indecisive’ region, all associated with low values of the bag parameter. The fig. 5 shows a clear trend of increasing evidence against EoSs with increasing values of the bag parameter. We found that 19 EoSs could be decisively rejected, whose bag parameter values were found to be $\geq$ 146.5. Furthermore, for $x_{v}=0.4$, 11 EoSs (with B $\geq$ 144.5) were found to lie in the ’decisive’ region, whereas, for $x_{v}=0.6$, 7 EoSs (with B $\geq$ 146.5) were found to lie in the ‘decisive’ region, implying that by increasing the value of the vector coupling constant, some higher values of the bag parameter could still be preferred by the current observations. The EoSs and their corresponding M-R curves for the rejected and accepted EoS are shown in fig. 6. Similar to the analysis of hadronic EoS, the SM EoSs also prefer stiffer EoS. This again shows that all the current astrophysical observations prefer a stiffer EoS. From fig. 7, we see the difference in MR curves when we do not include HESS J1731-347 data. Without the inclusion of this observation, only a few EoSs were found to be decisively rejected; however, this number significantly increased upon its inclusion.
III.3 Hybrid EoS
The hybrid EoSs constructed (a total of 637 are used in our analysis) in section II.3 were evaluated. XMLSLZ-GM1 with $x_{v}=0.5$ and B = 158.33 was found to have the highest evidence value. In the following sections, we shall denote this EoS as ‘Hyb_best’ to avoid lengthy phrases. The fig. 8 and fig. 14 (left) show the odds ratio plot of each EoS (for different $x_{v}$ values) with respect to Hyb_best.
A total of 429 EoSs were ‘decisively’ rejected when compared with Hyb_best, and there are 6 EoSs with odds ratio values situated in the ‘indecisive’ region. The comparable EoSs are: XMLSLZ-DDME2 ($x_{v}=0.5$, B=158.33), XMLSLZ-DDME2 ($x_{v}=0.44$, B=158.33), XMLSLZ-DDLZ1 ($x_{v}=0.44$, B=158.33), XMLSLZ-MTVTC ($x_{v}=0.44$, B=158.33), RG-SKb ($x_{v}=0.5$, B=155.0) and PT-GRDF2-DD2 ($x_{v}=0.5$, B=158.33). Additionally, it is observed that among the 7 comparable EoSs, including Hyb_best, 6 exhibit a bag parameter value of 158.33, indicating a greater preference for this specific value. However, it should be noted that such a preference is only observed in higher values of the vector coupling constant (fig. 8) and not in lower values. It is also evident in fig. 14 (left) that choosing B=158.33 for $x_{v}=0.2$ does not provide any improvement in the odds ratio value of the EoS family.
In fig. 8 and fig. 14 (left), there are certain EoS families that do not exhibit a lot of change in their evidence values, even after changing the vector coupling constant and bag parameter values. The fig. 9 shows the accepted and rejected hybrid EoSs (left) along with their corresponding M-R curves (right). The discontinuity in FOPT is represented in fig. 11 (left). This reveals that EoSs with early PT are preferred over those with PT at higher pressures.
The fig. 10 illustrates the M-R curves of hybrid EoSs when the data on HESS J1731-347 is not considered. Comparing fig. 9 (right) and fig. 10, we find that inclusion of the information of HESS J1731-347 changes the nature of EoSs which were rejected earlier. Without HESS J1731-347 information, only EoSs having higher radius were favored. Including the data of HESS J1731-347, stars with smaller radii cannot be rejected from our analysis. Analyzing fig. 11 (right) indicates that no comments can be made upon the transition of FOPT without the HESS data, but upon including it (left), one can definitely say that EoSs with early FOPTs are preferred by the observations.
Figure 8: Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.44; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.5. Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
Figure 9: Left: EoS plot of the hybrid EoSs; Right: Mass-Radius curves of the hybrid EoSs. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
Figure 10: Mass-Radius curves of the hybrid matter EoSs. The figure illustrates the rejected and accepted hybrid matter EoSs when HESS J1731-347 is not considered. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
Figure 11: Left: The density discontinuity corresponding to the hybrid EoSs from fig. 9 is shown; Right: The density discontinuity corresponding to the hybrid EoSs from fig. 10 is shown. The y-axis represents the pressure, and the x-axis the energy density. The color nomenclature remains the same as previous.
III.4 Comparison among the ‘Indecisive’ EoSs
Figure 12: Odds ratio plot of Hyb_best with the best performing EoSs in each family of EoSs. Note that EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
In the previous subsections, we have found the best performing EoS from the hadronic family, strange matter family, and the hybrid star family EoSs. The odds ratio analysis among the EoSs from each family that lie in the ‘indecisive’ region is performed. The fig. 12 shows the odds ratio plot of each EoS with Hyb_best. It shows that the hybrid EoS Hyb_best performs best among all the EoSs. We find that for the hybrid EoSs, there are no EoSs that lie beyond the ‘indecisive’ region, suggesting that the hybrid EoSs are the most probable among all the families of EoSs. The analysis of the hadronic EoSs reveals that EoSs following the RMF model, the Density-Dependent RMF model, and the Thomas-Fermi approach satisfy the current astrophysical observations the best. For the SM EoSs, we see that the EoSs having a value of B $\leq 139.5$ are most preferred.
IV Summary and Conclusions
The work revisits the current status of the EoSs upon including the mass and radius measurements of the compact object HESS J1731-347 using a bayesian model selection technique. The current astrophysical observations of - PSR J0030+0451, PSR J0740+6620, and GW170817, along with HESS J1731-347 were used to unravel the implications of the latter observation. Starting with constructing three families of EoSs - the first family consists of a hadronic matter comprised mostly of $npe\mu$. The second family of EoSs was built on the modified MIT bag model with vector couplings and comprising three flavored quarks. For the third family of EoSs, we used hybrid EoSs comprising of FOPT.
After obtaining the EoSs, a bayesian model selection was performed for each family of EoSs. In order to evaluate the odds ratio of the EoSs, the evidence integrals were solved by choosing mass as the parameter to be integrated. A uniform prior on mass was adopted, ranging from $0.5M_{\odot}$ to the maximum mass allowed by the EoS. While calculating the odds ratio, we assume each EoS is equally likely, thus reducing the odds ratio to the Bayes factor. The Jeffrey’s scale, which allows to decisively reject EoSs and choose the most suitable EoS models, was used.
The analysis of the hadronic family of EoSs shows that the nuclear EoSs following the Brueckner-Hartree-Fock approximation can be decisively rejected, and while the EoSs based on Effective interactions cannot be effectively rejected; however, model is one of the least effective models in explaining the current astrophysical observations along with the compact object HESS J1731-347. Thus, we conclude that if current observations are composed of hadronic matter, then they are unlikely to have an EoS based on these two models. SPG(M3) (unified NS EoS), following the density-dependent RMF model, was found to be the most likely. The density-dependent RMF model is one of the most effective model that satisfies all the current astrophysical observations best. None of the EoS models coming from the Brussels-Montreal energy density functionals could be decisively rejected. On further investigating the speed of sound of these EoSs, we found that at the intermediate density ranges, a stiffer EoS is preferred by the current astrophysical observations.
Analyzing the SM EoS, it was found that an increase in the Bag value fails to satisfy the current data. EoSs with bag parameter values greater than 146.5 were decisively rejected. Our analysis of the SM EoS shows that if current observations are SSs, then they would inherently prefer an EoS with smaller values of the bag parameter. However, we see that if we increase the value of the vector coupling constant, slightly higher bag values can still be preferred. The EoSs also show a similar trend to the hadronic EoSs, with the stiffer EoSs being preferred over the softer ones.
The analysis of the hybrid EoSs shows that a hybrid EoS referred to as Hyb_best was found to be the most likely hybrid EoS. It was found 6 EoSs comparable to Hyb_best, five of which also have bag parameter values equal to 158.33, indicating a greater preference towards this bag value. However, such a preference is only observed in higher values of the vector coupling constant $x_{v}$. Since the hybrid EoSs comparable to Hyb_best also had a value of B = 158.33, it suggests that contrary to our SM EoSs, including HESS J1731-347 in our analysis shows a preference for FOPT with larger Bag values. It was also found that the EoSs with FOPT at higher densities were likely disfavored compared to EoSs having an early PT.
Further comparison among all the families of EoSs that lie in the ‘indecisive’ region of our analysis, we find that all the hybrid EoSs perform the best, suggesting that explaining current astrophysical observations using a hybrid EoS is the most likely scenario. The SM EoSs also hint that the value of the bag parameter should be less than 139.5, such that all the astrophysical observations are satisfied best. We could not decisively comment on the hadronic EoS family. However, we found that some EoSs based on the RMF model, Density-Dependent RMF model, and Thomas-Fermi approximation are comparable to Hyb_best.
It is to be noted that since we have considered only the modified MIT Bag model with vector interactions as the quark counterpart in our hybrid EoSs, hence there also lies a possibility that the rejected hybrid EoSs might be able to explain the current observations upon considering different types of quark counterparts.
Acknowledgement
The authors would also like to thank IISER Bhopal for providing the infrastructure for this work. The authors also acknowledge Prasanta Char for their helpful discussions. DK acknowledges Bhaskar Biswas for email correspondence. ST would like to thank Akshat Singh, Pratik Thakur, Shamim Haque, and Rishi Gupta for their helpful discussions. SC acknowledges the Prime Minister’s Research Fellowship (PMRF), Ministry of Education Govt. of India, for a graduate fellowship. RM and DK acknowledge the Science and Engineering Research Board (SERB), Govt. of India, for monetary support in the form of a Core Research Grant (CRG/2022/000663).
Data Availability
The data can be availed upon reasonable request to the authors.
appendix
IV.1 Status of EoSs without HESS J1731-347
Figure 13: Left: Odds ratio plot of the hadronic matter EoS PCP(BSK25) with other hadronic matter EoSs. These are the results obtained when HESS J1731-347 was not taken into consideration. The color scheme is the same as in fig. 1; Right: Odds ratio plot of the strange matter EoS with $x_{v}=0.4$ and B$=139.0$ with other strange matter EoSs. These are the results obtained when HESS J1731-347 was not taken into consideration. The color scheme is the same as in fig. 1. The ticks on the y-axis refer to strange matter EoSs with the corresponding xv and B parameters.
In this section, we discuss the status of the EoSs in the context of astrophysical observations of only: PSR J0740+6620, PSR J0030+0451, and GW170817. Only including these observations, we illustrate how each family of hadronic, quark, and hybrid family of EoSs behave in fig. 13, fig. 14 (right), and fig. 15. In fig. 13 we show the odds ratio plot for the hadronic and SM EoSs. From the hadronic family of EoSs, we observe that none of the EoSs could be decisively rejected contrary to the results discussed upon including HESS observation. Furthermore, in this scenario, PCP(BSK25) (following the Brussels-Montreal energy density functional) was found to be the most likely EoS, with 29 other EoSs lying in the ’indecisive’ region. This is a significantly greater number than when the data of HESS is taken into consideration.
The SM EoS analysis shows that though there are several EoSs that can be decisively rejected based only on the other observations, however, the total number of EoSs that can be decisively rejected upon including HESS data significantly increases. In this scenario, the EoS with $x_{v}$ = 0.4 and B = 139.0 is found to be the most likely, with 17 EoSs lying in the ’indecisive’ region. Therefore, a similar trend is also followed by comparable equations, where the number of comparable equations decreases when the data of HESS J1731-347 is included.
Figure 14: Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.2 and when HESS was considered; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.2 and HESS was not considered. Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
[…]
Figure 15: The figure illustrates the odds ratio plot obtained when HESS J1731-347 was not considered. Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.44; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.5. Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.

### v2

@@
 Analyzing the hybrid family of equation of states consisting of a first-order phase transition revealed preferences for early first-order phase transition.
-Comparing all the preferred equations of state among each family it was found that the current astrophysical constraints most prefer the hybrid equation of states.
+Comparing all the preferred equations of state among each family, it was found that the current astrophysical constraints prefer the hybrid equation of states the most.
 preprint: APS/123-QED
@@
 The core of the NSs is still a mystery, with several models suggesting the presence of only hadronic matter [5, 6, 7].
-The possibility of PT at intermediate densities opens the possibility of hybrid stars [8, 9, 10, 11, 12], having an inner core composed of quark matter followed by an outer core of hadronic matter [13, 14, 15, 16, 17, 18].
+The possibility of PT at intermediate densities opens the possibility of hybrid stars [8, 9, 10, 11, 12, 13], having an inner core composed of quark matter followed by an outer core of hadronic matter [14, 15, 16, 17, 18, 19].
 There also exists another unique category of stars called strange stars (SSs) made up of strange quark matter (SQM) [20, 21, 22].
 Witten suggested that the SQM is the absolute ground state consisting of $u$,$d$, and $s$ quarks supporting the idea of SSs [23].
-The presence of SQM at the core of the star softens the equation of state (EoS) with possible violations in the maximum mass limit of the star imposed by PSR J0348+0432 [23] of mass $2.01\pm 0.04M_{\odot}$.
 The only way to demystify the core of NSs is with astrophysical observations.
@@
 With more improved observations in the future, the EoS band is expected to get thinner.
-However, the recent observation of HESS J1731-347 [34] started raising a few eyebrows with a mass and radius measurement of $0.77\pm^{0.20}_{0.17}M_{\odot}$ and $10.4\pm^{0.86}_{0.78}\text{km}$ respectively, making it the lightest known compact star till date.
-Since such low mass and radius are inherently favored by SSs, we cannot rule out the possibility of the idea of SQM inside the star.
-Several significant works followed this observation, each of them trying to explain the properties of this star [35, 36, 37, 38, 39].
-However, one of the best way to examine any observation is to do a statistical analysis with the data (basically different models) that one has.
+However, the recent observation of HESS J1731-347 [34] started raising a few eyebrows with a mass and radius measurement of $0.77\pm^{0.20}_{0.17}M_{\odot}$ and $10.4\pm^{0.86}_{0.78}\text{km}$ respectively, making it one of the lightest known compact stars till date.
+A recent work highlights that the analysis of the central compact object (CCO) in the supernova remnant (SNR) HESS J1731-347 is dependent on a number of necessary but insufficient conditions [35].
+In particular, the authors found that the assumption of a uniform temperature carbon atmosphere (UTCA) adopted by Ref.
+[34] is not at all consistent with the longest & highest quality XMM-Newton data, further stating that a UTCA model for the CCO is less likely.
+Although these points portray that the mass-radius calculation of the CCO is controversial, another recent work by Ref [36] talks about the minimal consistency checks for the observation of the remnant in HESS J1731-347 within existing models of SSs.
+They concluded with the remark that the CCO in the supernova remnant HESS J1731-347 passes initial consistency checks and can be utilised for theoretical and observational work.
+Furthermore, the recent work of Ref.
+[37] explores the possibility of HESS J1731-347 indeed being the lightest star observed.
+Their analysis found evidence that the star could indeed be a NS of mass 0.9 M formed from the collapse of a fast-rotating iron core.
+It is, therefore, necessary to at least check its effect on the present EoS bounds set by NS observations.
+Several significant works followed this observation, each of them trying to explain the properties of this star [38, 39, 40, 41, 42, 43, 44, 45].
+However, one of the best ways to examine any observation is to do a statistical analysis with the data (basically different models) one has.
 Bayesian analysis is an important tool that is being used significantly to constrain nuclear models [46, 47, 48].
 Recently, Ref [49] showed, using a Bayesian framework, the impact of the compact object in the context of nuclear matter.
-Bayesian inference techniques have been primarily used for constraining GW observation parameters [44, 45, 46] with Ref [47] developing a model selection technique using various astrophysical observations to compare nuclear matter EoSs.
-In this paper, we perform bayesian model selection to explore the implications of the compact object HESS J1731-347 along with other astrophysical observations in the light of a NS, a strange star, or a hybrid star.
-We consider several nuclear matter EoS models, models consisting of three flavored quarks, and also hybrid models based on first-order phase transition (FOPT).
+Bayesian inference techniques have been primarily used for constraining GW observation parameters [50, 51], with Ref [52] developing a model selection technique using various astrophysical observations to compare nuclear matter EoSs.
+Similarly, another paper presents the application of Bayesian model selection, ranking a few EoSs using only the GW data of GW170817 [53].
+By using only GW data, one can only constrain the equation of state at densities corresponding to the central density of canonical neutron star of mass 1.4M.
+In this paper, we perform Bayesian model selection to explore the implications of the compact object HESS J1731-347, along with other astrophysical observations in the light of a neutron star, a strange star, or a hybrid star.
+Since SSs inherently favor low mass and radius, we cannot rule out the possibility of such stars.
+Additionally, by using several astrophysical observations, we can constrain the EoS by surveying a wide range of densities.
+Furthermore, we consider several nuclear matter EoS models, models consisting of three flavored quarks, and hybrid models based on first-order phase transition (FOPT).
 Lastly, we compare the three sets of EoSs to analyze which family of EoSs explains the current astrophysical observations the best.
@@
 Section II discusses the formalism adopted in this work to describe the EoS models and also the model selection technique and how it is implemented in this work.
-The results are described in section III, and finally, important results and conclusion are summarized in section IV.
+The results are described in section III, and finally, in section IV, we conclude with a summary and discussion of our results.
 II Formalism
+In order to use the EoSs for astrophysical analysis, we first checked whether the EoSs satisfied constraints from chiral effective field theory.
+To do so, we adopted a construction similar to Refs.
+[33, 54], where, for densities in the range $0<n\leq 0.5n_{s}$, where $n_{s}$=0.16 fm-3, the Baym-Pethick-Sutherland (BPS) [55] EoS is considered.
+In the density range of $0.5n_{s}<n\leq 1.1n_{s}$, a series of monotropes of the form $p=K\rho^{\gamma}$ are constructed, where the value of ‘K’ is fixed by matching to the BPS EoS and $\gamma$ is varied in the range [1.77,3.23].
+These different monotropes in the density range $0.5n_{s}<n\leq 1.1n_{s}$ form our CET band.
+We then check whether all the EoSs that we use for our analysis lie within this region.
+Our analysis did not use any EoS that did not comply with this condition.
+In Appendix .1, we discuss the changes in our study when the CET constraint is not taken into consideration.
 II.1 Hadronic EoSs
-For the purpose of hadronic EoSs, we make use of the publicly available EoSs in the CompOSE repository [48].
-All these EoSs consist of $npe\mu$ nuclear matter.
+For the purpose of hadronic EoSs, we use the publicly available nuclear matter EoSs in the CompOSE repository [56].
 We have considered 7 EoS models, namely the Density-dependent Relativistic Mean Field (RMF) model, the Effective Interaction model, the Thomas Fermi approach, the RMF model, the Brussels-Montreal energy density functional, non-linear RMF model, and the Brueckner-Hartree-Fock approximation.
@@
 More details about the EoSs used can be found in table 1.
+Table 1: List of all the 38 hadronic EoS with the corresponding radius of a $0.77M_{\odot}$, $1.4M_{\odot}$, and $2M_{\odot}$ star, respectively; along with the Tidal Deformability measurement of the $1.4M_{\odot}$ star and the maximum mass given by each EoS.
+The last three equations are based on variational techniques, density-dependent covariant density functional, and generalized relativistic density functional, respectively.
+The information of the rejected hadronic EoSs is presented in bold text.
 EoS $R_{0.77}$(km) $R_{1.4}$(km) $R_{2.0}$(km) $\Lambda_{1.4}$ $M_{max}(\text{M}_{\odot})$
+Density Dependent RMF model
 GPPVA(TW)NSunifiedInnerCrust-core [57, 58] 12.75 12.33 11.42 401 2.07
@@
 SPG(M1)unifiedNSEoS [59] 12.78 12.8 12.87 534 2.54
+Effective Interactions
 RG(SkMp) [63, 64] 12.54 12.5 11.5 467 2.11
@@
 RG(SLY2) [63, 67] 11.93 11.79 10.7 307 2.06
-RG(SK255) [55, 60] 13.56 13.15 12.03 580 2.15
 RG(SLY230a) [63, 68] 11.90 11.83 11.05 324 2.11
 RG(SKa) [63, 66] 13.02 12.92 12.16 558 2.22
-RG(Rs) [55, 62] 13.01 12.93 11.84 553 2.12
 VGBCMR(D1MStar) [69] 11.67 11.71 10.47 314 2.00
-RG(SkI3) [55, 57] 13.53 13.55 12.80 771 2.25
-RG(SK272) [55, 60] 13.61 13.32 12.48 636 2.24
 RG(SLY9) [63, 67] 12.53 12.47 11.70 444 2.16
-RG(SKI2) [55, 57] 13.51 13.48 12.45 733 2.17
 RG(SkI6) [63, 65] 12.33 12.49 11.88 481 2.2
 RG(SLY4) [63, 70] 11.84 11.7 10.62 295 2.06
-RG(SkI5) [55, 57] 14.10 14.08 13.12 947 2.25
+Thomas Fermi approach
 XMLSLZ(DD-LZ1) [6, 71] 12.52 13.15 13.34 732 2.56
 XMLSLZ(DDME2) [6, 72] 12.74 13.2 13.22 712 2.48
-XMLSLZ(NL3) [6, 67] 14.30 14.59 14.60 1264 2.77
-XMLSLZ(PKDD) [6, 68] 13.75 13.63 13.14 765 2.33
-XMLSLZ(GM1) [6, 69] 13.47 13.76 13.38 892 2.36
-XMLSLZ(TM1) [6, 70] 14.31 14.28 13.51 1045 2.18
 XMLSLZ(DDME-X) [6, 73] 12.81 13.37 13.49 792 2.56
-XMLSLZ(MTVTC) [6, 72] 13.25 13.1 11.43 600 2.02
 XMLSLZ(TW99) [6, 58] 12.35 12.27 11.35 405 2.08
+RMF approximation
 PCGS(PCSB1) [74, 75] 12.98 13.25 12.67 624 2.19
@@
 PCP(BSK26) [59, 78] 11.7 11.77 11.18 323 2.17
+Brussels-Montreal energy density functionals
 PCP(BSK25) [59, 79] 11.97 12.37 12.10 476 2.22
@@
 PCP(BSK22) [59, 81] 12.97 13.04 12.58 624 2.26
+Nonlinear RMF models
 GPPVA(NL3wrL55)NSunifiedInnerCrust-core [82] 13.32 13.76 14.06 939 2.75
-GPPVA(FSU2)NSunifiedInnerCrust-core [83] 14.53 13.29 11.35 816 2.37
 GPPVA(FSU2H)NSunifiedInnerCrust-core [83] 12.91 13.29 10.26 750 2.37
 GPPVA(TM1e)NSunifiedInnerCrust-core [84] 13.02 13.16 10.59 661 2.12
+Brueckner-Hartree-Fock approximations
 BL(chiral)withUnifiedCrust [85] 12.60 12.27 11.13 386 2.08
@@
 PT(GRDF2-DD2)coldNS [88, 62] 12.84 13.17 13.07 686 2.42
-Table 1: List of all the 50 hadronic EoS with the corresponding radius of a $0.77M_{\odot}$, $1.4M_{\odot}$, and $2M_{\odot}$ star, respectively; along with the Tidal Deformability measurement of the $1.4M_{\odot}$ star and the maximum mass given by each EoS.
-EoSs in red are based on the density-dependent relativistic mean field (RMF) model, ones in crimson are based on Effective interactions, ones in brown are based on Thomas Fermi approach, ones in darkgreen are based on RMF approximations, ones in violet are based on Brussels-Montreal energy density functional, ones in magenta are based on non-linear RMF models and the ones in peach are based on Brueckner-Hartree-Fock approximation.
-The last three equations are based on variational techniques, density-dependent covariant density functional, and generalized relativistic density functional, respectively.
-The information of the rejected hadronic EoSs is presented with bold text.
 II.2 Quark Matter EoSs
@@
 II.3 EoSs with First-Order PT
-Assuming the phase transition to be of FOPT we construct the hybrid EoS from the hadronic and modified MIT bag-model quark matter EoS.
+Assuming the phase transition to be a FOPT, we construct the hybrid EoS from the hadronic and modified MIT bag-model quark matter EoS.
 The jump/transition from the hadronic phase (HP) to the quark phase (QP) happens at a particular pressure when the chemical potential of the quark phase becomes less than the chemical potential of the hadronic phase.
-Although this happens at a particular pressure and chemical potential, and they remain smooth throughout, there is a discontinuity in the energy density (and density) corresponding to the latent heat required for the transition.
-The transition is modeled using Maxwells construction as $p_{\rm HP}(\mu_{\rm c})=p_{\rm QP}(\mu_{\rm c})$ where $\mu_{\rm c}$ is the critical baryonic chemical potential where the transition occurs.
+Although this occurs at a specific pressure and chemical potential, and they remain smooth throughout, there is a discontinuity in the energy density (and density) corresponding to the latent heat required for the transition.
+First-order phase transition can be of two types depending on the surface tension between the adjoining fluids: Gibbs and Maxwell.
+The surface tension of the quark matter is the decisive parameter that dictates the type of phase transition construction mechanism.
+Although this parameter is poorly known, its theoretical estimates fall within a wide range of (5-300) MeV/fm3.
+In Refs.
+[92, 93] its value was found to lie in the range (10-50) MeV/fm3.
+However, in Ref.
+[94], authors estimated its value in the range (50-150) MeV/fm3 and in Ref.
+[95], an even higher value was estimated.
+Since we do not know the correct value of surface tension, both scenarios can be utilized to construct the phase transition.
+In the present study, we have considered Maxwell’s construction mechanism, with $p_{\rm HP}(\mu_{\rm c})=p_{\rm QP}(\mu_{\rm c})$ (where $\mu_{\rm c}$ is the critical baryonic chemical potential where the transition occurs), while considering a large value of surface tension.
+It should be noted that the FOPT could have also been modeled using Gibbs construction.
+However, this construction requires information on the chemical potential component-wise.
+Since CompOSE does not provide this information, we utilize the overall chemical potential of the state, provided by CompOSE, to perform Maxwell construction.
 II.4 Bayesian Model Selection
@@
 $P(M|d,I)=\dfrac{P(d|M,I)P(M|I)}{P(d|I)}$ (1)
-where $M$ refers to a model (EoS), $I$ refers to any background information that we have, and d refers to the astrophysical data.
+where $M$ refers to a model (EoS), $I$ refers to any background information we have, and d refers to the astrophysical data.
 $P(M|d,I)$ is the posterior probability of the model, $P(d|M,I)$ is the marginalized-likelihood (evidence) for the data, $P(M|I)$ is the prior probability and $P(d|I)$ is a constant term.
@@
 In order to compare two different models ($M_{1}$ and $M_{2}$), we find out the odds ratio between them, which is defined as:
-$\mathcal{O}_{M_{1}}^{M_{2}}=\dfrac{P(M_{2}|d,I)}{P(M_{1}|d,I)}=\dfrac{P(d|M_{2},I)}{P(d|M_{2},I)}\times\frac{P(M_{2}|I)}{P(M_{1}|I)}$ (3)
+$\mathcal{O}_{M_{1}}^{M_{2}}=\dfrac{P(M_{2}|d,I)}{P(M_{1}|d,I)}=\dfrac{P(d|M_{2},I)}{P(d|M_{1},I)}\times\frac{P(M_{2}|I)}{P(M_{1}|I)}$ (3)
+where the ratio of the likelihood for the two models is known as the Bayes factor.
 For uniformity, we can take the ratio of the priors to be equal to one so that $P(M_{2}|I)=P(M_{1}|I)$.
@@
 The mass and tidal deformability ($\Lambda$) measurements from GW170817 [29, 30, 96] serves as $d_{\rm GW}$.
-For $d_{\rm NICER}$, the mass and radius measurements from PSR J0030+0451 [27] and PSR J0740+6620 [25] serve as input.
+For $d_{\rm NICER}$, the mass and radius measurements from PSR J0030+0451 [27], PSR J0437-4715, and PSR J0740+6620 [25] serve as input.
 Similarly, the mass and radius measurements from HESS J1731-347 [34] form $d_{\rm HESS}$.
@@
 For every EoS, we solve the Tolman-Oppenheimer-Volkoff (TOV) equations [97] to obtain the mass-radius curve, also known as the MR curve.
-Our EoS can be parametrized either by using the mass or radius values obtained after solving the TOV equations.
-In our scenario, we use the mass values, and hence the evidence for the NICER observations is given by:
-$P(d_{\rm NICER}|EoS,I)$ $=\>\int_{m_{min}}^{m_{max}}P(d_{\rm NICER}|m,$
+Our EoS can be parametrized either by using the mass or by the radius values obtained after solving the TOV equations.
+In our scenario, we use the mass values, and hence, the evidence for the NICER observations is given by:
+$P(d_{\rm NICER}|{\rm EoS},I)$ $=\>\int_{m_{\rm min}}^{m_{\rm max}}P(d_{\rm NICER}|m,$
 $R(m,EoS),EoS,I)$
@@
 Similarly, for the HESS observation, the evidence is given as :
-$P(d_{\rm HESS}|EoS,I)$ $=\>\int_{m_{min}}^{m_{max}}P(d_{\rm HESS}|m,$
+$P(d_{\rm HESS}|EoS,I)$ $=\>\int_{m_{\rm min}}^{m_{\rm max}}P(d_{\rm HESS}|m,$
 $R(m,EoS),EoS,I)$
@@
 Without loss of generality, we can choose a uniform prior on mass [52, 49]. It is given by:
-$\begin{split}P(m|EoS,I)&=\frac{1}{m_{max}-m_{min}};\>\text{$m_{min}\leq m\leq m_{max}$}\\ &=0\hskip 62.59596pt;\text{everywhere else}\end{split}$ (9)
+$\begin{split}P(m|EoS,I)&=\frac{1}{m_{\rm max}-m_{\rm min}};\>\text{$m_{\rm min}\leq m\leq m_{\rm max}$}\\ &=0\hskip 62.59596pt;\text{everywhere else}\end{split}$ (9)
 $m_{max}$ is the maximum mass of the EoS obtained after solving the TOV equation.
@@
 To calculate the evidence for the GW data, we parameterize the two masses of binaries ($m_{1}$,$m_{2}$) and their corresponding tidal deformabilities ($\lambda_{1}$,$\lambda_{2}$) as:
-$P(d_{\rm GW}|EoS,I)$ $=\int_{m_{2}}^{M_{max}}dm_{1}\int_{M_{min}}^{m_{1}}P(d_{\rm GW}|m_{1},m_{2},$
+$P(d_{\rm GW}|EoS,I)$ $=\int_{m_{2}}^{M_{\rm max}}dm_{1}\int_{M_{\rm min}}^{m_{1}}P(d_{\rm GW}|m_{1},m_{2},$
 $\lambda_{1}(EoS,m_{1}),\lambda_{2}(EoS,m_{2}),EoS,I)$
@@
 To solve eq (10) we make use of the chirp mass [98] given by:
-$\mathcal{M}_{chirp}=\dfrac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}=1.186M_{\odot}$ (11)
+$\mathcal{M}_{\rm chirp}=\dfrac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}=1.186M_{\odot}$ (11)
 Where $m_{1}$ and $m_{2}$ are the masses of the primary and secondary neutron stars having a mass ratio, $q={m_{2}}/{m_{1}}\geq 0.73$ inferred from GW170817 observation [96].
@@
 We construct the likelihood using a multivariate Gaussian KDE with the mass and tidal deformability samples from the observation.
-All of the evidence integrals were performed using PyMultinest, which is a Python package for implementing the MultiNest algorithm.
+All of the evidence integrals were performed using PyMultiNest [99], which is a Python package for implementing the MultiNest algorithm.
 It offers efficient evidence calculation for multi-modal data.
-Furthermore, all likelihood distributions were constructed using the multivariate KDE method of Statsmodels.
+Furthermore, all likelihood distributions were constructed using the multivariate KDE method of Statsmodels [100].
 III Results
-For our analysis, the observations we have used are $(i)$ GW170817, $(ii)$ Three X-Ray sources, namely PSR J0030+0451[27], PSR J0740+6620 [25], and HESS J1731-347 [34].
-For the three sets of EoSs that we have considered in our analysis, we have 50 hadronic EoSs, 58 strange matter EoSs, and 637 hybrid EoSs.
+For our analysis, the observations we have used are $(i)$ GW170817, $(ii)$ Three X-Ray sources, namely PSR J0030+0451[27], PSR J0437-4715[101], PSR J0740+6620 [25], and HESS J1731-347 [34].
+Appendix .2 outlines the results of our analysis when the data of HESS J1731-347 is not taken into consideration.
+For the three sets of EoSs we have considered in our analysis, we have 38 hadronic EoSs, 58 strange matter EoSs, and 544 hybrid EoSs.
 Since the hybrid EoSs were constructed using Maxwell construction, for each nuclear EoS considered, we agnostically generated a family of hybrid EoSs, resulting in a large number of hybrid EoSs.
 We adopt Jeffrey’s [102] scale for the log of the odds ratio values.
-Note that Jeffrey’s scale is defined for the Bayes Factor and not the odds ratio, but upon taking the ratio of the prior of each model to be unity, the odds ratio becomes equal to the Bayes Factor.
+Jeffrey’s scale is defined for the Bayes factor, not the odds ratio.
+However, upon taking the ratio of the prior of each model to be unity, the odds ratio becomes equal to the Bayes factor.
 According to the scale, if $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-0.5, 0), then although there is evidence for model $M_{1}$, it is not worth more than a bare mention.
@@
 If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ is smaller than -2, then there is ‘decisive’ evidence for (against) model $M_{1}$ ($M_{2}$), and we can reject model $M_{2}$.
-Utilizing this, we present our analysis in the next subsections.
+Utilizing this, we present our analysis in the following subsections.
 III.1 Hadronic EoS
-Figure 1: Odds ratio plot of SPG(M3) with other hadronic EoSs.
+Figure 1: Odds ratio plot of XMLSLZ(DD-LZ1) with other hadronic EoSs.
 Following Jeffrey’s scale, the region between $(-0.5,0)$ (shaded with ‘plum’) is the region in which, if a model lies, it provides evidence for the base model, but it is not worth more than a bare mention (indecisive).
@@
 The region beyond $-2$ (shaded with ‘misty rose’) is the region of decisive evidence for the base model.
-The histogram for the odds ratio value of an EoS is depicted in the corresponding color of the respective region it lies in; for example, if an EoS is situated in the decisive region, its histogram is colored in ‘mistyrose’.
-After evaluating the evidence value for each EoS, the EoS with the highest evidence value is SPG(M3) (with unified NS EoS), which uses density-dependent relativistic mean-field approximation.
+The histogram for the odds ratio value of an EoS is depicted in the corresponding colour of the respective region it lies in; for example, if an EoS is situated in the decisive region, its histogram is coloured in ‘misty rose’.
+$EoS_{i}$ refers to the equation of state being compared.
+After evaluating the evidence value for each EoS, the EoS with the highest evidence value is XMLSLZ(DD-LZ1), which uses the Thomas Fermi approach.
 Fig.
-1 shows the odds ratio plot of each EoS with respect to SPG(M3).
-Upon utilizing Jeffrey’s scale, as discussed before, there are exactly 23 EoS that can be decisively rejected based on their odds ratio value with respect to SPG(M3).
+1 shows the odds ratio plot of each EoS with respect to XMLSLZ(DD-LZ1).
+Upon utilizing Jeffrey’s scale, as discussed before, there are precisely 15 EoS that can be decisively rejected based on their odds ratio value with respect to XMLSLZ(DD-LZ1).
 Table 1 highlights the rejected EoSs with bold text.
-Additionally, there are 9 EoSs with odds ratio values situated in the ‘indecisive’ region (see fig.
-1 for reference).
+Additionally, there are 9 other EoSs with odds ratio values situated in the ‘indecisive’ region (fig.
+1).
 They are based on the following EoS models: density-dependent RMF model, Thomas Fermi approximation, RMF model, and generalized relativistic density functional.
-Figure 2: Left: EoS plot of the hadronic EoSs.
-The grey patch shows the region where we see a change in the stiffness of EoS.
-The fig.
-3 shows the speed of analysis in the region of grey patch.
-Right: Mass-Radius curves of the hadronic EoSs.
-The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
+Figure 2: (a): EoS plot of the hadronic EoSs.
+The grey patch shows the region where there is a change in the stiffness of EoS; (b): Mass-Radius curves of the hadronic EoSs; (c): Illustrates the speed of sound corresponding to the shaded region in (a); (d): Illustrates the rejected and accepted hadronic EoSs when HESS J1731-347 is not considered.
+In all the figures, rejected EoSs are plotted in dark blue, while the accepted ones are plotted in salmon.
 Using table 1 as a reference, it can be observed that except the EoSs based on Brussels-Montreal energy density functional, every other type of EoS has at least one EoS that is being rejected.
 However, none of the EoSs based on the Brussels-Montreal energy density functional lie in the ‘indecisive’ region.
-On the contrary, while one EoS based on the density-dependent RMF model can be decisively rejected, 5 (out of 7) of the EoSs lie in the ‘indecisive’ region, with SPG(M3) as the best performing EoS.
-This analysis suggests that the density-dependent RMF model best satisfies all the current astrophysical observations.
+On the contrary, while one of the EoS based on the density-dependent RMF model can be decisively rejected, 5 (out of 7) of the EoSs lie in the ‘indecisive’ region.
+Similarly, 3 (out of 4) of the EoSs following the Thomas Fermi approximation lie in the ‘indecisive’ region, with XMLSLZ(DD-LZ1) as the best performing EoS.
+This analysis suggests that the density-dependent RMF model and the Thomas Fermi approximation best satisfy all the current astrophysical observations.
+Additionally, the EoS PT(GRDF2-DD2) (based on generalized relativistic density functional) also performed exceptionally well.
+Using table 1 and fig.
+1, it can be seen that the effective-interaction based model, although not decisively rejected, is the least effective model when explaining current astrophysical bounds.
 Furthermore, it should be noted that all EoSs based on the Brueckner-Hartree-Fock approximation were rejected, leading us to the conclusion that it is the least plausible EoS type that could explain the current observations, including HESS J1731-347.
-It was found that 12 (out of 16) Effective interactions based EoSs could be decisively rejected.
-Among the remaining four EoSs, none are situated in the comparable region; hence, among the EoS types that could not be decisively rejected, the Effective interaction type EoS is the least effective model in our analysis.
-Furthermore, APR(APR) (based on variational techniques) lies in the ‘strong’ region, CMGO(GDFM-1) (based on density-dependent covariant density functional) lies in the ‘strong’ region, and PT(GRDF2-DD2) (based on generalized relativistic density functional) lies in the ‘substantial’ region.
-In fig.
-2, we show the comparison of the EoSs that were rejected along with their mass-radius curves.
-We see that at intermediate density ranges, the EoSs which are softer (having less speed of sound) are rejected compared to the stiffer EoSs, which is also shown in fig.
-3.
-The adiabatic speed of sound ($c_{s}=\sqrt{\partial p/\partial\epsilon}$) is an important quantity as it determines the slope of the EoS.
-A stiffer EoS would tend to explain higher-mass stars.
+Most of the EoSs were accepted or rejected based on either the radius or the tidal deformability bound.
+Since the observations in consideration are rather recent, those EoSs that were old and did not maintain the radius or tidal deformability bounds were found to be rejected.
+Additionally, using table 1 and fig.
+1 as a reference, it can be seen that the comparable EoSs have very similar radii for R1.4 and R2.0.
+We notice a trend of increasing evidence against EoSs that have increasingly dissimilar values for R1.4 and R2.0.
+A similar trend was also observed by Ref.
+[6].
+In figs.
+2(a) and 2(b), we show the comparison of the EoSs that were rejected along with their MR curves.
 Fig.
-3 shows an increasing behaviour in the speed of sound which is characteristic of hadronic matter EoSs [97, 98, 99, 100, 101].
-Contrary to the scenario when we do not consider the observation of HESS J1731-347 none of the hadronic EoSs used in this work could be decisively rejected as shown in fig.
-4.
+2(b) shows that the accepted EoSs have an MR curve that has a back-bending effect above 0.5 M.
+They also support a larger maximum mass.
+Additionally, the back-bending of the MR curve is reflected in the sudden stiffening of the EoS beyond $200$ MeVfm-3 in fig.
+2(a).
+However, at much higher densities, the curves become softer.
+This shows that the best performing EoSs are more non-monotonous than the rejected ones.
+This non-monotonous nature can be examined in greater detail in the speed of sound plot.
+The adiabatic speed of sound ($c_{s}=\sqrt{\partial p/\partial\epsilon}$) is an essential quantity as it determines the slope of the EoS [103, 104, 105, 106, 107].
+Fig.
+2(c) shows an interesting feature; the accepted curves are clearly more non-monotonic than the rejected curves, with a few even attaining a local maximum.
+Usually, this is associated with the production of certain new degrees of freedom or suppression of some existing degrees [108, 109, 110].
+Fig.
+2(d) shows the rejected and accepted mass-radius curves of the hadronic EoSs when the data of HESS J1731-347 was not considered.
+Contrary to the scenario when we consider the observation of HESS J1731-347, none of the hadronic EoSs used in this work could be decisively rejected in this scenario, as shown in fig.
+2(d).
 The rejection of the EoSs from our analysis shows that all the current astrophysical observations prefer a stiffer hadronic EoS at intermediate densities.
-Figure 3: The speed of sound corresponding to the shaded region in fig.
-2 (left) is shown here.
-The y-axis represents the speed of sound in geometrized units, and the x-axis represents the energy density.
-The color nomenclature remains the same as previous plots.
-Figure 4: Mass-Radius curves of the hadronic matter EoSs.
-The figure illustrates the rejected and accepted hadronic EoSs when HESS J1731-347 is not considered.
-Since none of the EoSs are rejected, all EoSs are plotted in salmon.
 III.2 Strange EoS
-Figure 5: Odds ratio plot of the strange matter EoS with $x_{v}=0.6$ and B = 139.5 with other strange matter EoSs.
-The color scheme is the same as in fig.
+Figure 3: Odds ratio plot of the strange matter EoS with $x_{v}=0.4$ and B = 139.0 with other strange matter EoSs.
+The colour scheme is the same as in fig.
 1.
 The ticks on the y-axis refer to strange matter EoSs with the corresponding $x_{v}$ and B parameters.
-The strange matter EoSs were constructed by varying the vector coupling constant ‘$x_{v}$’ and bag parameters ‘B’ by small intervals as discussed in section II.2.
-Out of the 58 constructed EoSs, the EoS with $x_{v}=0.6$ and B=139.5 possesses the highest evidence value.
-Figure 6: Left: EoS plot of the strange matter EoSs; Right: Mass-Radius curves of the strange matter EoSs. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
-Figure 7: Mass-Radius curves of the strange matter EoSs.
-The figure illustrates the rejected and accepted strange matter EoSs when HESS J1731-347 is not considered.
-The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
-The fig.
-5 shows the odds ratio plot of each quark matter EoS in comparison to the EoS with $x_{v}=0.6$ and B = 139.5.
-There are 12 EoSs situated in the ‘indecisive’ region, all associated with low values of the bag parameter.
-The fig.
-5 shows a clear trend of increasing evidence against EoSs with increasing values of the bag parameter.
-We found that 19 EoSs could be decisively rejected, whose bag parameter values were found to be $\geq$ 146.5.
-Furthermore, for $x_{v}=0.4$, 11 EoSs (with B $\geq$ 144.5) were found to lie in the ’decisive’ region, whereas, for $x_{v}=0.6$, 7 EoSs (with B $\geq$ 146.5) were found to lie in the ‘decisive’ region, implying that by increasing the value of the vector coupling constant, some higher values of the bag parameter could still be preferred by the current observations.
-The EoSs and their corresponding M-R curves for the rejected and accepted EoS are shown in fig.
-6.
-Similar to the analysis of hadronic EoS, the SM EoSs also prefer stiffer EoS.
-This again shows that all the current astrophysical observations prefer a stiffer EoS.
-From fig.
-7, we see the difference in MR curves when we do not include HESS J1731-347 data.
+$EoS_{i}$ refers to the equation of state being compared.
+The strange matter EoSs were constructed by varying the scaled coupling constant ‘$x_{v}$’ and bag parameters ‘B’ by small intervals as discussed in section II.2.
+Out of the 58 constructed EoSs, the EoS with $x_{v}=0.4$ and B=139.0 possesses the highest evidence value.
+Figure 4: (a): EoS plot of the strange matter EoSs; (b): Mass-Radius curves of the strange matter EoSs; (c): Mass-Radius curves of the strange matter EoSs when HESS J1731-347 is not considered.
+The colour nomenclature remains the same as the previous one.
+Fig.
+3 shows the odds ratio plot of each quark matter EoS in comparison to the EoS with $x_{v}=0.4$ and B = 139.0.
+There are 12 other EoSs situated in the ‘indecisive’ region, all associated with low values of the bag parameter.
+Fig.
+3 shows a clear trend of increasing evidence against EoSs with increasing values of the bag parameter, with only two minor deviations from it.
+This should be attributed to the fact that increasing the value of the bag parameter decreases the stiffness of the EoS.
+Therefore, our analysis suggests that similar to the hadronic matter EoSs, strange matter EoSs also prefer stiff EoSs.
+We found that 17 EoSs could be decisively rejected, whose bag parameter values were found to be $\geq$ 146.5.
+Furthermore, for $x_{v}=0.4$, 10 EoSs (with B $\geq$ 145.0) were found to lie in the ’decisive’ region, whereas, for $x_{v}=0.6$, 7 EoSs (with B $\geq$ 146.5) were found to lie in the ‘decisive’ region, implying that by increasing the value of the scaled coupling constant, some higher values of the bag parameter could still be preferred by the current observations.
+This is an expected result since an increase in the value of the scaled coupling constant increases the stiffness of the EoS.
+The rejected and accepted EoSs and their corresponding M-R curves are shown in figs.
+4(a) and 4(b) respectively.
+In the M-R plot, the EoSs that barely cut the PSR J0030+0451 contour are rejected, and those that have a significant overlap with the contour are accepted by our analysis.
+Looking at fig.
+4(a), it can be seen that at low pressure, there is a cutoff value of density for the EoS to be rejected (approximately 225 MeVfm-3).
+From figs.
+4(b) and 4(c), we see the difference in MR curves when we do not include HESS J1731-347 data.
 Without the inclusion of this observation, only a few EoSs were found to be decisively rejected; however, this number significantly increased upon its inclusion.
 III.3 Hybrid EoS
-The hybrid EoSs constructed (a total of 637 are used in our analysis) in section II.3 were evaluated.
-XMLSLZ-GM1 with $x_{v}=0.5$ and B = 158.33 was found to have the highest evidence value.
+The hybrid EoSs constructed (545 are used in our analysis) in section II.3 were evaluated.
+XMLSLZ-DDME2 with $x_{v}=0.44$ and B = 158.33 was found to have the highest evidence value.
 In the following sections, we shall denote this EoS as ‘Hyb_best’ to avoid lengthy phrases.
-The fig.
-8 and fig.
-14 (left) show the odds ratio plot of each EoS (for different $x_{v}$ values) with respect to Hyb_best.
-A total of 429 EoSs were ‘decisively’ rejected when compared with Hyb_best, and there are 6 EoSs with odds ratio values situated in the ‘indecisive’ region.
-The comparable EoSs are: XMLSLZ-DDME2 ($x_{v}=0.5$, B=158.33), XMLSLZ-DDME2 ($x_{v}=0.44$, B=158.33), XMLSLZ-DDLZ1 ($x_{v}=0.44$, B=158.33), XMLSLZ-MTVTC ($x_{v}=0.44$, B=158.33), RG-SKb ($x_{v}=0.5$, B=155.0) and PT-GRDF2-DD2 ($x_{v}=0.5$, B=158.33).
-Additionally, it is observed that among the 7 comparable EoSs, including Hyb_best, 6 exhibit a bag parameter value of 158.33, indicating a greater preference for this specific value.
-However, it should be noted that such a preference is only observed in higher values of the vector coupling constant (fig.
-8) and not in lower values.
-It is also evident in fig.
-14 (left) that choosing B=158.33 for $x_{v}=0.2$ does not provide any improvement in the odds ratio value of the EoS family.
-In fig.
-8 and fig.
-14 (left), there are certain EoS families that do not exhibit a lot of change in their evidence values, even after changing the vector coupling constant and bag parameter values.
-The fig.
-9 shows the accepted and rejected hybrid EoSs (left) along with their corresponding M-R curves (right).
-The discontinuity in FOPT is represented in fig.
-11 (left).
-This reveals that EoSs with early PT are preferred over those with PT at higher pressures.
-The fig.
-10 illustrates the M-R curves of hybrid EoSs when the data on HESS J1731-347 is not considered.
+Figs.
+5 and LABEL:fig:Hybrid_xv02 show the odds ratio plot of each EoS (for different $x_{v}$ values) with respect to Hyb_best.
+A total of 351 EoSs were ‘decisively’ rejected when compared with Hyb_best, and there are 4 other EoSs with odds ratio values situated in the ‘indecisive’ region.
+The comparable EoSs are: XMLSLZ-DDME2 ($x_{v}=0.5$, B=158.33), XMLSLZ-DDLZ1 ($x_{v}=0.44$, B=158.33), RG-SKb ($x_{v}=0.5$, B=155.0) and PT-GRDF2-DD2 ($x_{v}=0.5$, B=158.33).
+Additionally, it is observed that among the five comparable EoSs, including Hyb_best, four of them exhibit a bag parameter value of 158.33, indicating a greater preference for this specific value.
+However, it should be noted that such a preference is only observed in higher values of the scaled coupling constant (fig.
+5) and not in lower values.
+It is also evident in LABEL:fig:Hybrid_xv02 that choosing B=158.33 for $x_{v}=0.2$ does not improve the odds ratio value of the EoS family.
+In figs.
+5 and LABEL:fig:Hybrid_xv02, certain EoS families do not exhibit much change in their evidence values, even after changing the scaled coupling constant and bag parameter values.
+Figs.
+7(a) and 7(b) show the accepted and rejected hybrid EoSs and their corresponding M-R curves, respectively.
+Since it is difficult to comment on the hybrid EoSs using only fig.
+7(b), we utilize the two parameters of Maxwell’s construction: width of the discontinuity and the onset of phase transition.
+Fig.
+8(a) illustrates both these parameters for the rejected and accepted EoSs.
+It can be seen that EoSs with an early onset of PT are preferred over those with a PT at higher densities.
+We conclude that the onset density of PT is more important, and the preference of the EoS does not depend on the width of the discontinuity.
+Fig.
+7(c) illustrates the M-R curves of hybrid EoSs when the data on HESS J1731-347 is not considered.
 Comparing fig.
-9 (right) and fig.
-10, we find that inclusion of the information of HESS J1731-347 changes the nature of EoSs which were rejected earlier.
-Without HESS J1731-347 information, only EoSs having higher radius were favored.
+7(b) and fig.
+7(c), we find that the inclusion of the information of HESS J1731-347 dramatically increases the number of EoSs being rejected and puts a better constraint on the nature of the EoS.
+Without HESS J1731-347 information, only EoSs having higher radii were favoured.
 Including the data of HESS J1731-347, stars with smaller radii cannot be rejected from our analysis.
 Analyzing fig.
-11 (right) indicates that no comments can be made upon the transition of FOPT without the HESS data, but upon including it (left), one can definitely say that EoSs with early FOPTs are preferred by the observations.
-Figure 8: Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.44; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.5.
-Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
-Figure 9: Left: EoS plot of the hybrid EoSs; Right: Mass-Radius curves of the hybrid EoSs. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
-Figure 10: Mass-Radius curves of the hybrid matter EoSs.
-The figure illustrates the rejected and accepted hybrid matter EoSs when HESS J1731-347 is not considered.
-The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
-Figure 11: Left: The density discontinuity corresponding to the hybrid EoSs from fig.
-9 is shown; Right: The density discontinuity corresponding to the hybrid EoSs from fig.
-10 is shown.
-The y-axis represents the pressure, and the x-axis the energy density.
-The color nomenclature remains the same as previous.
+8(b) indicates that no comments can be made upon the transition of FOPT without the HESS data.
+However, upon including it (fig.
+8(a)), one can definitely say that EoSs with early FOPTs are preferred by the observations.
 III.4 Comparison among the ‘Indecisive’ EoSs
-Figure 12: Odds ratio plot of Hyb_best with the best performing EoSs in each family of EoSs. Note that EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
-In the previous subsections, we have found the best performing EoS from the hadronic family, strange matter family, and the hybrid star family EoSs.
-The odds ratio analysis among the EoSs from each family that lie in the ‘indecisive’ region is performed.
-The fig.
+In the previous subsections, we have found the best performing EoS from the hadronic family, strange matter family, and the hybrid star family of EoSs.
+In this section, the odds ratio analysis among the EoSs from each family that lie in the ‘indecisive’ region is performed.
+Fig.
 9 shows the odds ratio plot of each EoS with Hyb_best.
 It shows that the hybrid EoS Hyb_best performs best among all the EoSs.
-We find that for the hybrid EoSs, there are no EoSs that lie beyond the ‘indecisive’ region, suggesting that the hybrid EoSs are the most probable among all the families of EoSs.
+We find that for the hybrid EoSs, no EoSs lie beyond the ‘indecisive’ region, suggesting that the hybrid EoSs are the most probable among all the families of EoSs.
 The analysis of the hadronic EoSs reveals that EoSs following the RMF model, the Density-Dependent RMF model, and the Thomas-Fermi approach satisfy the current astrophysical observations the best.
 For the SM EoSs, we see that the EoSs having a value of B $\leq 139.5$ are most preferred.
+Figure 5: (a): Odds ratio plot of Hyb_best with EoSs having a fixed $x_{v}$ value of 0.44; (b): Odds ratio plot of Hyb_best with EoSs having a fixed $x_{v}$ value of 0.5.
+The colour nomenclature of the disks is the same as in figs.
+1 and 3.
+Each sector belongs to a hadronic family.
+The different shaped points correspond to different bag values referenced in (a).
+The different colours of the points help distinguish continuous sets of hadronic families.
+The odds ratio value of each EoS acts as the radius value in the plot.
+This value is then plotted in the sector of the hadronic family of that EoS, with the shape of its corresponding bag value.
+Figure 7: (a): EoS plot of the hybrid EoSs; (b): Mass-Radius curves of the hybrid EoSs; (c): Mass-Radius curves of the hybrid EoSs when HESS J1731-347 is excluded.
+The colour nomenclature remains the same as the previous one.
+Figure 8: (a): The density discontinuity corresponding to the hybrid EoSs from fig.
+7(b) is shown; (b): The density discontinuity corresponding to the hybrid EoSs from fig.
+7(c) is shown.
+The colour nomenclature remains the same as the previous one.
+Figure 9: Odds ratio plot of Hyb_best with the best performing EoSs in each family of EoSs. Note that EoS above refers to Hyb_best, and $EoS_{i}$ refers to the equation of state being compared.
 IV Summary and Conclusions
-The work revisits the current status of the EoSs upon including the mass and radius measurements of the compact object HESS J1731-347 using a bayesian model selection technique.
-The current astrophysical observations of - PSR J0030+0451, PSR J0740+6620, and GW170817, along with HESS J1731-347 were used to unravel the implications of the latter observation.
-Starting with constructing three families of EoSs - the first family consists of a hadronic matter comprised mostly of $npe\mu$.
-The second family of EoSs was built on the modified MIT bag model with vector couplings and comprising three flavored quarks.
-For the third family of EoSs, we used hybrid EoSs comprising of FOPT.
+The work improves the current status of the EoSs upon including the mass and radius measurements of the compact object HESS J1731-347 using a Bayesian model selection technique.
+The current astrophysical observations of - PSR J0030+0451, PSR J0740+6620, J0437-4715, and GW170817, along with HESS J1731-347, were used to unravel the implications of the latter observation.
+Starting with constructing three families of EoSs - the first family consists of hadronic matter.
+The second family of EoSs was built on the modified MIT bag model with scaled couplings and comprising of three flavoured quarks.
+For the third family of EoSs, we used hybrid EoSs comprising a FOPT.
 After obtaining the EoSs, a Bayesian model selection was performed for each family of EoSs.
@@
 While calculating the odds ratio, we assume each EoS is equally likely, thus reducing the odds ratio to the Bayes factor.
-The Jeffrey’s scale, which allows to decisively reject EoSs and choose the most suitable EoS models, was used.
-The analysis of the hadronic family of EoSs shows that the nuclear EoSs following the Brueckner-Hartree-Fock approximation can be decisively rejected, and while the EoSs based on Effective interactions cannot be effectively rejected; however, model is one of the least effective models in explaining the current astrophysical observations along with the compact object HESS J1731-347.
-Thus, we conclude that if current observations are composed of hadronic matter, then they are unlikely to have an EoS based on these two models.
-SPG(M3) (unified NS EoS), following the density-dependent RMF model, was found to be the most likely.
-The density-dependent RMF model is one of the most effective model that satisfies all the current astrophysical observations best.
-None of the EoS models coming from the Brussels-Montreal energy density functionals could be decisively rejected.
-On further investigating the speed of sound of these EoSs, we found that at the intermediate density ranges, a stiffer EoS is preferred by the current astrophysical observations.
-Analyzing the SM EoS, it was found that an increase in the Bag value fails to satisfy the current data.
-EoSs with bag parameter values greater than 146.5 were decisively rejected.
-Our analysis of the SM EoS shows that if current observations are SSs, then they would inherently prefer an EoS with smaller values of the bag parameter.
-However, we see that if we increase the value of the vector coupling constant, slightly higher bag values can still be preferred.
-The EoSs also show a similar trend to the hadronic EoSs, with the stiffer EoSs being preferred over the softer ones.
-The analysis of the hybrid EoSs shows that a hybrid EoS referred to as Hyb_best was found to be the most likely hybrid EoS.
-It was found 6 EoSs comparable to Hyb_best, five of which also have bag parameter values equal to 158.33, indicating a greater preference towards this bag value.
-However, such a preference is only observed in higher values of the vector coupling constant $x_{v}$.
-Since the hybrid EoSs comparable to Hyb_best also had a value of B = 158.33, it suggests that contrary to our SM EoSs, including HESS J1731-347 in our analysis shows a preference for FOPT with larger Bag values.
-It was also found that the EoSs with FOPT at higher densities were likely disfavored compared to EoSs having an early PT.
-Further comparison among all the families of EoSs that lie in the ‘indecisive’ region of our analysis, we find that all the hybrid EoSs perform the best, suggesting that explaining current astrophysical observations using a hybrid EoS is the most likely scenario.
-The SM EoSs also hint that the value of the bag parameter should be less than 139.5, such that all the astrophysical observations are satisfied best.
-We could not decisively comment on the hadronic EoS family.
-However, we found that some EoSs based on the RMF model, Density-Dependent RMF model, and Thomas-Fermi approximation are comparable to Hyb_best.
-It is to be noted that since we have considered only the modified MIT Bag model with vector interactions as the quark counterpart in our hybrid EoSs, hence there also lies a possibility that the rejected hybrid EoSs might be able to explain the current observations upon considering different types of quark counterparts.
+We adopted Jeffrey’s scale, allowing us to decisively reject EoSs and choose the most suitable EoS models.
+The analysis of the hadronic family of EoSs shows that the nuclear EoSs following the Brueckner-Hartree-Fock approximation and EoSs based on effective interactions are least effective in explaining the current astrophysical observations along with the compact object HESS J1731-347.
+XMLSLZ(DD-LZ1), following the Thomas Fermi approximation, was found to be the best performing EoS.
+Density-dependent RMF models and the Thomas Fermi approximation are the most effective models in explaining the current observations.
+Most of the EoSs were accepted or rejected based on either the radius or the tidal deformability bound of the observations.
+The accepted EoSs showed a back-bending effect on the MR curve, which is due to a sudden stiffening of the EoS.
+Moreover, these models showed a non-monotonous nature in their speed of sound, with few showing maxima, which can be associated with the emergence of new degrees of freedom at higher densities.
+Analyzing the SM EoS, it was found that an increase in the Bag value fails to satisfy the current data as it effectively makes the EoS soft.
+Our analysis of the SM EoS shows that if current observations are SSs, they would inherently prefer an EoS with smaller values of the bag parameter.
+However, if we increase the value of the scaled coupling constant, slightly higher bag values can still be preferred.
+There is a clear distinction between the rejected and accepted SM EoSs, both in the EoS and MR plot, with the data of the pulsar PSR J0030+0451 playing a decisive role in the segregation of the SM EoSs.
+Additionally, the analysis hints towards an energy density cut-off value at low pressures.
+If the energy density of an EoS at low pressure lies above this threshold value, it is rejected by our analysis.
+The analysis of the hybrid EoSs shows that a hybrid EoS referred to as Hyb_best is the most likely hybrid EoS having a bag parameter value equal to 158.33 and scaled coupling constant value equal to 0.44.
+It was found that 4 other EoSs are comparable to Hyb_best, three of which also have bag parameter values equal to 158.33, indicating a greater preference towards this bag value.
+However, such a preference is only observed in higher values of the scaled coupling constant $x_{v}$.
+Although the MR curve does help in distinguishing the accepted EoSs from the rejected EoSs, this segregation is more dependent on the onset density of phase transition, preferring early PT densities.
+On comparing all the EoSs of different families that lie in the ‘indecisive’ region of our analysis, we find that all the hybrid EoSs perform the best, suggesting that explaining current astrophysical observations using a hybrid EoS is the most likely scenario.
+The results are consistent with the recent observation that EoSs show a non-monotonic speed of sound.
+The non-monotonicity is usually associated with the appearance of extra degrees of freedom (in this case, quark matter), suggesting a phase transition at intermediate densities in neutron stars.
+This is reflected in Hyb_best being the most favored EoS.
+It is to be noted that we have considered only the modified MIT Bag model with vector interactions as the quark counterpart in our hybrid EoSs, hence there lies a possibility that the rejected hybrid EoSs might be able to explain the current observations upon considering different types of quark counterparts.
+Since the current analysis does not consider the Gibbs construction, there also lies a possibility of explaining the present astrophysical observations using a FOPT constructed using the Gibbs construction.
+All such extensions are our future endeavors.
 Acknowledgement
@@
 RM and DK acknowledge the Science and Engineering Research Board (SERB), Govt.
-of India, for monetary support in the form of a Core Research Grant (CRG/2022/000663).
+of India, for monetary support through a Core Research Grant (CRG/2022/000663).
 Data Availability
-The data can be availed upon reasonable request to the authors.
-appendix
-IV.1 Status of EoSs without HESS J1731-347
-Figure 13: Left: Odds ratio plot of the hadronic matter EoS PCP(BSK25) with other hadronic matter EoSs.
+The data can be made available upon reasonable request by the authors.
+Appendix A:   Effect of the CET Constraint
+In section II, we discussed the construction of the CET band in order to use it to filter EoSs.
+This section discusses the changes in our analysis when the CET constraint is not considered.
+Initially, our analysis included 50 hadronic EoSs and 637 hybrid EoSs.
+We found that from the 50 hadronic EoSs considered, 12 of them did not satisfy the CET bound.
+On performing the entire analysis with these EoSs, we saw that 8 out of the 12 EoSs that did not satisfy the CET constraint were already rejected from our analysis.
+Of the remaining 4 EoSs, two were in the ‘substantial’ regime, and two were in the ‘strong’ regime.
+Similarly, on performing the analysis for the initial 637 hybrid EoSs, we found that 93 EoSs did not satisfy the CET constraint.
+Out of these 92 EoSs, 86 were rejected from our analysis, 2 were in the ‘strong’ regime, 3 were in the ‘substantial’ regime, and 1 was found to be in the ‘indecisive’ regime.
+Appendix B:   Status of EoSs without HESS J1731-347
+Figure 10: (a): Odds ratio plot of the hadronic matter EoS PT(GRDF2-DD2) with other hadronic matter EoSs.
 These are the results obtained when HESS J1731-347 was not taken into consideration.
-The color scheme is the same as in fig.
-1; Right: Odds ratio plot of the strange matter EoS with $x_{v}=0.4$ and B$=139.0$ with other strange matter EoSs.
+The colour scheme is the same as in fig.
+1; (b): Odds ratio plot of the strange matter EoS with $x_{v}=0.4$ and B$=139.0$ with other strange matter EoSs.
 These are the results obtained when HESS J1731-347 was not taken into consideration.
-The color scheme is the same as in fig.
+The colour scheme is the same as in fig.
 1.
 The ticks on the y-axis refer to strange matter EoSs with the corresponding xv and B parameters.
-In this section, we discuss the status of the EoSs in the context of astrophysical observations of only: PSR J0740+6620, PSR J0030+0451, and GW170817.
-Only including these observations, we illustrate how each family of hadronic, quark, and hybrid family of EoSs behave in fig.
-13, fig.
-14 (right), and fig.
-15.
+In this section, we discuss the status of the EoSs in the context of astrophysical observations of only PSR J0740+6620, PSR J0030+0451, PSR J0437-4715, and GW170817.
+Only including these observations, we illustrate how each of the hadronic, quark, and hybrid family of EoSs behave in figs.
+10, LABEL:fig:Hybrid_pulsarplot and LABEL:fig:Hybrid_xv02_onlypulsar.
 In fig.
-13 we show the odds ratio plot for the hadronic and SM EoSs.
+10, we show the odds ratio plot for the hadronic and SM EoSs.
 From the hadronic family of EoSs, we observe that none of the EoSs could be decisively rejected contrary to the results discussed upon including HESS observation.
-Furthermore, in this scenario, PCP(BSK25) (following the Brussels-Montreal energy density functional) was found to be the most likely EoS, with 29 other EoSs lying in the ’indecisive’ region.
-This is a significantly greater number than when the data of HESS is taken into consideration.
-The SM EoS analysis shows that though there are several EoSs that can be decisively rejected based only on the other observations, however, the total number of EoSs that can be decisively rejected upon including HESS data significantly increases.
-In this scenario, the EoS with $x_{v}$ = 0.4 and B = 139.0 is found to be the most likely, with 17 EoSs lying in the ’indecisive’ region.
+Furthermore, in this scenario, PT(GRDF2-DD2) (following the Brussels-Montreal energy density functional) was the most likely EoS, with 24 other EoSs lying in the ’indecisive’ region.
+This is a significantly greater number than when the data of HESS is considered.
+The SM EoS analysis shows that although several EoSs can be decisively rejected based only on these observations, the total number of EoSs that can be decisively rejected upon including HESS data significantly increases.
+In this scenario, the EoS with $x_{v}$ = 0.4 and B = 139.0 is the most likely, with 15 other EoSs lying in the ’indecisive’ region.
 Therefore, a similar trend is also followed by comparable equations, where the number of comparable equations decreases when the data of HESS J1731-347 is included.
-Figure 14: Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.2 and when HESS was considered; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.2 and HESS was not considered.
-Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
 […]
-Figure 15: The figure illustrates the odds ratio plot obtained when HESS J1731-347 was not considered.
-Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.44; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.5.
-Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.